# Complexity of equal 0-surgeries

### v2

@@ 2. The low-crossing knots
In this section, we work with the low-crossing knots, i.e. the prime knots that have diagrams of at most $15$ crossings. We start with the classification of their 00-surgeries.
@@
Tables 2 and 3 list the Regina names of all non-hyperbolic 00-surgeries of prime knots with at most $15$ crossings. Any knot that is not contained in that table is proven to have a hyperbolic 00-surgery. The list of the hyperbolic volumes of these 00-surgeries can be accessed at [AKW24].
The complete list of prime, non-hyperbolic knots of at most $15$ crossings and their descriptions as torus knots or satellite knots is given in Table 4. Any 00-surgery on such a knot is again non-hyperbolic.
@@
SnapPy has implemented a verified version of the volume function, that if successful, returns an interval of small length in which the volume provable has to lie. In that case, the manifold is also proven to be hyperbolic. If unsuccessful an error is returned. In that case, the manifold might be hyperbolic or not.
We use that verified volume function in SnapPy to compute the volume of the 00-surgeries. If SnapPy fails to compute the verified volume, we try to find a better triangulation where it works or raise the precision for the computations. If both methods fail to compute the volume, we use the Regina recognition code [Dun20a], cf. [FPS24] to recognize the triangulation of the 00-surgery as a triangulation of a non-hyperbolic manifold. This successfully classifies all 00-surgeries of the low-crossing knots. ∎
Table 2. This table collects all non-hyperbolic 00-surgeries along knots with at most $15$ crossings, together with their Regina names. Here $A$ stands for an annulus, $P$ for a pair of pants, and $M$ for a Möbius strip.
@@
$K13n3523$ $SFS\,[A:(2,1)]/[2,7|1,3]$
@@
Table 3. Table 2 continued.
DT name non-hyperbolic name
@@
Next, we create a census of all friends $(K,K^{\prime})$ such that $K$ and $K^{\prime}$ are both prime and have at most $15$ crossings. It is known that knots with diffeomorphic 00-surgeries share the same Alexander polynomial. So we first run through the list of all prime knots with at most $15$ crossings, compute their Alexander polynomials, and sort the knots into groups that have the same Alexander polynomials. If it turns out that some group contains only a single knot $K$, then this knot cannot have a friend that is prime and that has at most $15$ crossings, and we do not consider $K$ anymore. (The possibility that $K$ and $-K$ are not-isotopic but still are friends is discussed in the proof of Proposition 4.3 below.)
So in the following, we only consider groups of knots with the same Alexander polynomials that contain more than one knot. In the next step, we use the results from Theorem 2.1 to further split up the groups into knots that have all non-hyperbolic or all hyperbolic 00-surgeries. If the 00-surgery is hyperbolic, we also split into groups of knots where the 00-surgeries have overlapping volumes. By using the verified symmetry group computations in SnapPy we can further distinguish a few more 00-surgeries.
Some of the non-hyperbolic 00-fillings we can distinguish by their Regina names. For example, using the pieces in their JSJ decomposition and the classification of Seifert fibered spaces. Note, however, that the gluing map of the JSJ pieces is not returned. So if two Regina names agree or disagree, that does not necessarily mean that the manifolds are diffeomorphic or non-diffeomorphic.
Next, we know that several other knot invariants agree for knots that share the same 00-surgeries. If two knots share the same 00-surgery, they have the same $3$-genus, the same fiberedness status [Gab87], the same signatures, and isomorphic knot Floer homologies groups in certain gradings [OS04, Corollary 4.5], cf. [BS22a, Remark 2.2]. We use Szabó’s knot Floer homology calculator [Sza] to compute these invariants and refine our groups further.
For the remaining groups of knots that might share the same 00-surgeries, we try to distinguish their 00-surgeries by showing that their fundamental groups are not isomorphic, as done for example in [Bur20, Dun20b]. If $H$ is a subgroup of a fundamental group $G$ of a 00-surgery and $N$ the core of $H$, then the tuple
@@
is an invariant of the subgroup $H>G$. We use SnapPy, Regina, and GAP [GAP22] to determine all conjugacy classes of subgroups of $G$ up to index $7$ and compute for each the above invariants. If the collections of these invariants do not agree, then the fundamental groups are not isomorphic and thus the 00-surgeries are not diffeomorphic.
This was enough to distinguish the 00-surgery of most pairs of low-crossing knots. However, for around $200$ pairs of knots these invariants either all agreed or were too time-consuming to compute. The remaining pairs we could distinguish using their length spectra. The verified computations of the length spectra in SnapPy up to length $4$ were enough to distinguish all but $6$ pairs.
Among the remaining $6$ pairs of knots, we check if the 00-surgeries are diffeomorphic. For hyperbolic 00-surgeries this is done with SnapPy by searching for an isometry. For the non-hyperbolic 00-surgeries we load the triangulations to Regina and search through the Pachner graphs for a combinatorial equivalence between the triangulations.
Note that the SnapPy search for isometries between hyperbolic $3$-manifold also returns orientation-reversing isometries. But we want to have orientation-preserving isometries. For that, we use the following method to search for orientation-preserving isometries. For closed manifolds, SnapPy does not give any information on whether the isometry is orientation-preserving or not. To determine if two closed $3$-manifolds $M_{1}$ and $M_{2}$ are orientation preserving-diffeomorphic, we can use SnapPy to drill out curves $c_{1}$ and $c_{2}$ from $M_{1}$ and $M_{2}$. For cusped manifolds, SnapPy also displays the action of the isometry on the cusp from which one can read-off whether the isometry is orientation-preserving or not. Then we can search for an orientation-preserving isometry from $M_{1}\setminus c_{1}$ to $M_{2}\setminus c_{2}$ that fixes the meridians of $c_{1}$ and $c_{2}$. This isometry extends to an orientation-preserving isometry from $M_{1}$ to $M_{2}$.
In total, we find $6$ pairs of friends among the low-crossing knots (and below two more pairs, each consisting of a knot and its mirror). These are displayed in Table 1. All other pairs of prime knots with at most $15$ crossings have non-diffeomorphic 00-surgeries.
@@ 3. Connected sums
Next, we show that no knot among the low crossing knots has 00-surgery equal to the 00-surgery of a connected sum. For that, we first prove the following lemma.
@@
Let $K=K_{1}\#\cdots\#K_{n}$ be a connected sum of non-trivial, non-satellite knots $K_{1},\ldots,K_{n}$. If $K^{\prime}$ is another non-trivial connected sum such that $K$ and $K^{\prime}$ have orientation-preserving diffeomorphic 00-surgeries. Then $K$ and $K^{\prime}$ are isotopic.
@@
It is well-known (see for example [Bud06]) that the exterior $E_{K}$ of $K$ decomposes as
@@
Now it follows that performing the 00-surgery on $K$ is the same as capping off the outer boundary component of $D_{n}$ to get a $2$-sphere $S_{n}$ with $n$ holes. Thus the JSJ decomposition of the 00-surgery of $K$ is given as
@@
Now let $K^{\prime}=K^{\prime}_{1}\#\cdots\#K^{\prime}_{n^{\prime}}$ be another connected sum of non-trivial knots $K^{\prime}_{1},\ldots,K^{\prime}_{n^{\prime}}$ such that $K(0)$ is orientation-preserving diffeomorphic to $K^{\prime}(0)$. Then the uniqueness of the JSJ decomposition implies that the JSJ decompositions of $K(0)$ and $K^{\prime}(0)$ are the same. In particular, it follows that $n=n^{\prime}$ and that (after renumbering) $E_{K_{i}}$ is orientation-preserving diffeomorphic to $E_{K^{\prime}_{i}}$. Then [GL89] tells us that $K_{i}$ is isotopic to $K^{\prime}_{i}$.
@@
With that lemma, we are ready to prove the following proposition which allows us to restrict to prime knots.
@@
From the proof of Lemma 3.1 we know that the 00-surgery of a connected sum decomposes as
$(S_{n}\times S^{1})\cup\bigcup_{i=1}^{n}E_{K_{i}}.$
However, from Tables 2 and 3 we see directly that there are only two 00-surgeries that might be of that form: The 00-surgeries of $K_{+}=K15n142188$ and $K_{-}=K15n156076$ are both given by gluing the exterior $E_{K3a1}$ of the trefoil knot $K3a1$ to a $1$-cusped hyperbolic manifold $M_{\pm}$. To finish the proof, it is enough to show that in both cases the manifold $M_{\pm}$ is not the exterior of a knot in $S^{3}$.
From Table 4 we see that $K_{\pm}$ are satellites of the trefoil $K3a1$ with pattern ±$\pm L6a1$. It follows that $M_{\pm}$ is given by $w_{\pm}$-surgery on one component [^3] Since $L6a1$ has an isometry that interchanges the components (and preserve the framing) it does not matter which component we fill. of ±$\pm L6a1$, where $w_{\pm}=12\pm 3$ is the writhe of the diagrams of $K_{\pm}$ shown in Figure 1. From that description, we can build the manifold $M_{\pm}$ in SnapPy, compute the short slopes of $M_{\pm}$, and verify that the corresponding fillings are not homology spheres. Then the $6$-theorem [Ago00, Lac00] implies that $M_{\pm}$ is not the exterior of a knot in $S^{3}$.
For the census knots, all exceptional fillings are classified by Dunfield [Dun20a]. From his list, we can read-off as above that no 00-surgery can be a connected sum. ∎
@@
$K14n22180$ $(-3)$-fold twisted positive Whitehead double of $K3a1$
$K14n26039$ $(-3)$-fold twisted negative Whitehead double of $K3a1$
@@
$K15n59184$ $(-2)$-fold twisted negative Whitehead double of $K3a1$
$K15n115646$ $(-4)$-fold twisted positive Whitehead double of $K3a1$
@@
$K15n142188$ $(-3)$-fold twisted satellite of $K3a1$ with pattern $L6a1$, see Figure 1
$K15n156076$ $(-3)$-fold twisted satellite of $K3a1$ with pattern $-L6a1$, see Figure 1
Figure 1. Diagrams of $K15n142188$ (left) and $K15n156076$ (right) showing their satellite structures on $K3a1$.
@@ 4. Mirrors
  • The mirror $-K_{m}$ of $K_{m}$ is isotopic to $K_{-m}$,
  • $K_{m}$ is not isotopic to its mirror $-K_{m}$, but
@@
We will describe the infinite family via an RGB link construction as for example in [MP18, Pic19, Pic20, MP23, BKM25]. For that we consider the $3$-component hyperbolic link $RGB$ with components $R$, $G$, and $B$ in Figure 2(i). By rotating that diagram around the $z$-axis, we see that there exists an involution $f\colon S^{3}\setminus RGB\rightarrow S^{3}\setminus RGB$, that maps $R$ to itself but interchanges $G$ and $B$.
@@
Now we see $RGB$ as a Kirby diagram by dotting $R$ and adding framings $-m-1$ and $m-1$ to $G$ and $B$, respectively. Note that $R$ is a meridian of both $G$ and $B$ (but not simultaneously). Thus we can cancel $R$ and $G$ and denote the image of $B$ under this cancellation by $K_{B}^{-m}$. By symmetry, we can also reverse the roles of $B$ and $G$ and get a knot $K_{G}^{m}$. By construction, these knots share the same trace. From the above symmetry of the RGB link, it follows that $K_{B}^{m}$ is isotopic to $K_{G}^{m}$ and we define this knot to be $K_{m}$. Thus we have constructed knots $K_{m}$ such that $K_{m}$ and $K_{-m}$ share the same trace. Figure 2(ii)-(iii) show explicit Kirby moves to create out of the RGB link the knots $K_{m}$.
@@
Next we show that $-K_{m}=K_{-m}$ and that the $K_{m}$ are mutually distinct. Since the $K_{m}$ are obtained from $K_{0}$ by twisting three strands $m$ times, it follows from [KMS92] that infinitely many of the $K_{m}$ are distinct. More concretely, we can see that they are all mutually distinct. For that, we consider the surgery description of the $K_{m}$ along the $2$-component link $L$ shown in Figure 2(iv). With SnapPy we check that $L$ is a hyperbolic link. By computing the volumes of the $K_{m}$ for small values of $m$ and using [NZ85] we conclude that the volumes of the $K_{m}$ converge monotonically to the volume of $L$ and in particular all $K_{m}$ (for $m\in\mathbb{N}$) have different volumes.
@@
Moreover, we use SnapPy to compute the symmetry group of $L$, from which we observe that $L$ is isotopic to its mirror. This implies that $K_{-m}$ is isotopic to the mirror $-K_{m}$ of $K_{m}$.
@@
It remains to demonstrate the each of the $K_{m}$ is not isotopic to its mirror. This can also be deduced from the symmetry group of $L$. Using SnapPy, we check that the complement of $L$ admits no non-trivial orientation-preserving isometry that extends to the link. Thus Thursthon’s hyperbolic Dehn filling theorem [Thu79] implies that any sufficiently large Dehn filling of $L$ yields manifolds without any isometry, see [BKM24b] for details. The $K_{m}$ arise by Dehn filling one component of $L$ with slope $1/m$. Since the length of $1/m$ converges monotonically to $\infty$ it follows that infinitely many of the $K_{m}$ are asymmetric and thus not isotopic to their mirrors. More concretely, we can show that in fact for all $m\in\mathbb{N}$ the knots $K_{m}$ are not isotopic to their mirrors. For that we use the quantification of Thurston’s hyperbolic Dehn filling theorem [FPS22] (see for example [BKM24b]) to compute an explicit bound $B$ such that $K_{m}$ is asymmetric for all $m\geq B$. For the finitely remaining $m$, we use SnapPy to create triangulations of the complements of $K_{m}$ and check that their symmetry groups do not contain any orientation-reversing diffeomorphisms. (But $K_{1}$ and $K_{2}$ turn out to be not asymmetric.) ∎
$-m-1$
$-m-1$
$K_{G}^{-m}=K_{-m}$
$m-1$
$m-1$
00
@@
Let $K$ be a prime knot with at most $15$ crossings or a hyperbolic knot with tetrahedral complexity at most $9$. Then $K$ and $-K$ have orientation-preserving diffeomorphic 00-surgeries if and only if $K$ is amphicheiral or if $K$ is isotopic to one of the following six chiral knots: $K13n469=v2272$, $K14n3411$, $K15n64176$, $t11462$, $o9\_22951$, and $o9\_41058$.
@@
If $K$ and $K^{\prime}$ have orientation-preserving diffeomorphic 00-surgeries then their signatures agree $\sigma(K)=\sigma(K^{\prime})$. For any knot $K$ we have $\sigma(K)=-\sigma(-K)$. Thus if $K$ and $-K$ have orientation-preserving diffeomorphic 00-surgeries, then the signature of $K$ has to vanish.

### v3

@@ 2. The low-crossing knots
In this section, we work with the low-crossing knots, i.e. the prime knots that have diagrams of at most $15$ crossings. We start with the classification of their 0-surgeries.
@@
Tables 2 and 3 list the Regina names of all non-hyperbolic 0-surgeries of prime knots with at most $15$ crossings. Any knot that is not contained in that table is proven to have a hyperbolic 0-surgery. The list of the hyperbolic volumes of these 0-surgeries can be accessed at [3].
The complete list of prime, non-hyperbolic knots of at most $15$ crossings and their descriptions as torus knots or satellite knots is given in Table 4. Any 0-surgery on such a knot is again non-hyperbolic.
@@
SnapPy has implemented a verified version of the volume function, that if successful, returns an interval of small length with rational boundaries that is guaranteed to contain the exact value of the hyperbolic volume. In that case, the manifold is also proven to be hyperbolic. If unsuccessful, an error is returned. In that case, the manifold might be hyperbolic or not.
We use that verified volume function in SnapPy to compute for each 0-surgery a small interval in which the exact value of the volume lies. If SnapPy fails to compute the verified volume, we try to find a better triangulation where it works or raise the precision for the computations. If both methods fail to compute the volume, we use the Regina recognition code [22], cf. [26] to recognize the triangulation of the 0-surgery as a triangulation of a non-hyperbolic manifold. This successfully classifies all 0-surgeries of the low-crossing knots. ∎
Table 2. This table collects all non-hyperbolic 0-surgeries along knots with at most $15$ crossings, together with their Regina names [22]. Here AA stands for an annulus, $P$ for a pair of pants, and $M$ for a Möbius strip. If $X$ is a census manifold with 22-boundary components, then $JSJ(X)$ denotes a closed manifold obtained by gluing the boundaries of $X$ together with some diffeomorphism.
@@
$K13n3523$ SFS[A:(2,1)]/[2,7|1,3]SFS\,[A:(2,1)]/[2,7|1,3]
Table 3. Table 2 continued.
DT name non-hyperbolic name
@@
Next, we create a census of all friends $(K,K^{\prime})$ such that $K$ and $K^{\prime}$ are both prime and have at most $15$ crossings. It is known that knots with diffeomorphic 0-surgeries share the same Alexander polynomial. So we first run through the list of all prime knots with at most $15$ crossings, compute their Alexander polynomials, and sort the knots into groups that have the same Alexander polynomials. If it turns out that some group contains only a single knot $K$, then this knot cannot have a friend that is prime and that has at most $15$ crossings, and we do not consider $K$ anymore. (The possibility that $K$ and K-K are not-isotopic but still are friends is discussed in the proof of Proposition 4.3 below.)
So in the following, we only consider groups of knots with the same Alexander polynomials that contain more than one knot. In the next step, we use the results from Theorem 2.1 to further split up the groups into knots that have all non-hyperbolic or all hyperbolic 0-surgeries. If the 0-surgery is hyperbolic, we also split into groups of knots where the intervals, containing the volumes of the 0-surgeries, overlap. By using the verified symmetry group computations in SnapPy we can further distinguish a few more 0-surgeries.
Some of the non-hyperbolic 0-surgeries we can distinguish by their Regina names [22]. For example, using the pieces in their JSJ decomposition and the classification of Seifert fibered spaces. Note, however, that the gluing map of the JSJ pieces is not returned. So if two Regina names agree or disagree, that does not necessarily mean that the manifolds are diffeomorphic or non-diffeomorphic.
Next, we know that several other knot invariants agree for knots that share the same 0-surgeries. If two knots share the same 0-surgery, they have the same $3$-genus, the same fiberedness status [27], the same signatures, and isomorphic knot Floer homologies groups in certain gradings [44, Corollary 4.5], cf. [14, Remark 2.2]. We use Szabó’s knot Floer homology calculator [50] to compute these invariants and refine our groups further.
For the remaining groups of knots that might share the same 0-surgeries, we try to distinguish their 0-surgeries by showing that their fundamental groups are not isomorphic, as done for example in [17, 23]. If $H$ is a subgroup of a fundamental group $G$ of a 0-surgery and $N$ the core of $H$, then the tuple
@@
is an invariant of the subgroup $H<G$. We use SnapPy, Regina, and GAP [28] to determine all conjugacy classes of subgroups of $G$ up to index $7$ and compute for each the above invariants. If the collections of these invariants do not agree, then the fundamental groups are not isomorphic and thus the 0-surgeries are not diffeomorphic.
This was enough to distinguish the 0-surgery of most pairs of low-crossing knots. However, for around $200$ pairs of knots these invariants either all agreed or were too time-consuming to compute. The remaining pairs we could distinguish using their length spectra. The verified computations of the length spectra in SnapPy up to length $4$ were enough to distinguish all but $6$ pairs.
Among the remaining $6$ pairs of knots, we check if the 0-surgeries are diffeomorphic. For hyperbolic 0-surgeries this is done with SnapPy by searching for an isometry. For the non-hyperbolic 0-surgeries we load the triangulations to Regina and search through the Pachner graphs for a combinatorial equivalence between the triangulations.
Note that the SnapPy search for isometries between hyperbolic $3$-manifold also returns orientation-reversing isometries. But we want to have orientation-preserving isometries. For that, we use the following method to search for orientation-preserving isometries. For closed manifolds, SnapPy does not give any information on whether the isometry is orientation-preserving or not. To determine if two closed $3$-manifolds $M_{1}$ and $M_{2}$ are orientation preserving-diffeomorphic, we can use SnapPy to drill out curves $c_{1}$ and $c_{2}$ from $M_{1}$ and $M_{2}$. For cusped manifolds, SnapPy also displays the action of the isometry on the cusp from which one can read off whether the isometry is orientation-preserving or not. Then we can search for an orientation-preserving isometry from $M_{1}\setminus c_{1}$ to $M_{2}\setminus c_{2}$ that fixes the meridians of $c_{1}$ and $c_{2}$. Such an isometry extends uniquely over the Dehn fillings to an orientation-preserving isometry $M_{1}\to M_{2}$, and by construction this extension maps the core curve $c_{1}$ to $c_{2}$.
In total, we find $6$ pairs of friends among the low-crossing knots (and below two more pairs, each consisting of a knot and its mirror). These are displayed in Table 1. All other pairs of prime knots with at most $15$ crossings have non-diffeomorphic 0-surgeries.
@@ 3. Connected sums
Next, we show that no knot among the low crossing knots has 0-surgery equal to the 0-surgery of a connected sum. For that, we first prove the following general lemma.
@@
Let $K=K_{1}\#\cdots\#K_{n}$ be a connected sum of non-trivial, non-satellite knots $K_{1},\ldots,K_{n}$. If $K^{\prime}$ is another non-trivial connected sum such that $K$ and $K^{\prime}$ have orientation-preserving diffeomorphic 0-surgeries, then $K$ and $K^{\prime}$ are isotopic.
@@
It is well-known (see for example [16]) that the exterior $E_{K}$ of $K$ decomposes as
@@
Now it follows that performing the 0-surgery on $K$ is the same as capping off the outer boundary component of $D_{n}$ to get a $2$-sphere $S_{n}$ with $n$ holes. Thus the JSJ decomposition of the 0-surgery of $K$ is given as
@@
Now let $K^{\prime}=K^{\prime}_{1}\#\cdots\#K^{\prime}_{n^{\prime}}$ be another connected sum of non-trivial, prime knots $K^{\prime}_{1},\ldots,K^{\prime}_{n^{\prime}}$ such that $K(0)$ is orientation-preserving diffeomorphic to $K^{\prime}(0)$. Then the uniqueness of the JSJ decomposition implies that the JSJ decompositions of $K(0)$ and $K^{\prime}(0)$ are the same, i.e.
$K(0)=(S_{n}\times S^{1})\cup\bigcup_{i=1}^{n}E_{K_{i}}=K^{\prime}(0).$
On the other hand, we can decompose $K^{\prime}(0)$ as
$K^{\prime}(0)=(S_{n^{\prime}}\times S^{1})\cup\bigcup_{i=1}^{\,n^{\prime}}E_{K^{\prime}_{i}}.$
Note that the latter might not be the JSJ decomposition since $E_{K^{\prime}_{i}}$ might have a non-trivial JSJ decomposition (this happens if $K^{\prime}_{i}$ is a satellite knot).
Next, we compare these two decompositions to show that $n=n^{\prime}$ and (after renumbering) $E_{K_{i}}$ is orientation-preserving diffeomorphic to $E_{K^{\prime}_{i}}$.
By assumption, each $E_{K_{i}}$ is either the exterior of a hyperbolic knot or a torus knot. In particular, no $E_{K_{i}}$ is a product. If $n=2$, then $K(0)$ contains no product and thus also $n^{\prime}=2$. It follows that (after renumbering) $E_{K_{i}}$ is orientation-preserving diffeomorphic to $E_{K^{\prime}_{i}}$.
If $n>2$, then the JSJ decomposition of $K(0)=K^{\prime}(0)$ contains a unique piece of the form $S_{n}\times S^{1}$, and thus we have two cases. The first case is that $n^{\prime}=2$ and exactly one of the $E_{K^{\prime}_{i}}$ contains a JSJ piece of the form $S_{n}\times S^{1}$. The second case is $n=n^{\prime}$, then $S_{n}\times S^{1}$ is mapped to $S_{n^{\prime}}\times S^{1}$ and (after renumbering) $E_{K_{i}}$ is orientation-preserving diffeomorphic to $E_{K^{\prime}_{i}}$. We argue that the first case cannot occur. Indeed, suppose that $n^{\prime}=2$ and that one of the exteriors $E_{K^{\prime}_{i}}$ contains a JSJ piece diffeomorphic to $S_{n}\times S^{1}$. Since $K^{\prime}_{i}$ is assumed to be prime, it has to be a satellite knot. Then, by the JSJ classification of knot exteriors due to Budney [16], the JSJ decomposition of $E_{K^{\prime}_{i}}$ contains a piece that is the exterior of a link in $S^{3}$ with at least two components. On the other hand, since none of the knots $K_{i}$ is a satellite by assumption, the JSJ decomposition of $K(0)$ contains no JSJ piece other than $S_{n}\times S^{1}$ which has more than one boundary component. This contradicts the uniqueness of the JSJ decomposition. Hence, the first case cannot occur.
In summary we have $n=n^{\prime}$ and (after renumbering) $E_{K_{i}}$ is orientation-preserving diffeomorphic to $E_{K^{\prime}_{i}}$. Then [29] tells us that $K_{i}$ is isotopic to $K^{\prime}_{i}$.
@@
From the proof of the above lemma, we get, in particular, the following decomposition theorem for the 0-surgery of a connected sum.
Corollary 3.2
Let $K=K_{1}\#\cdots\#K_{n}$ be a connected sum of non-trivial knots $K_{i}$, then the 0-surgery $K(0)$ decomposes as
$K(0)=(S_{n}\times S^{1})\cup\bigcup_{i=1}^{n}E_{K_{i}},$
where $E_{K_{i}}$ is the exterior of $K_{i}$, $S_{n}$ denotes a $2$-sphere with $n$ holes, and where the gluing is along incompressible tori.
In particular, for $n=2$, it follows that the 0-surgery is obtained by gluing the exteriors of $K_{1}$ and $K_{2}$, i.e. $K(0)=E_{K_{1}}\cup E_{K_{2}}$.∎
With that corollary, we are ready to prove the following proposition which allows us to restrict to prime knots.
@@
From Corollary 3.2 we know that the 0-surgery of a non-trivial connected sum decomposes along incompressible tori as
$(S_{n}\times S^{1})\cup\bigcup_{i=1}^{n}E_{K_{i}}$
or for $n=2$ as $E_{K_{1}}\cup E_{K_{2}}$. However, from Tables 2 and 3 we see directly that there are only two 0-surgeries that might be of that form: The 0-surgeries of $K_{+}=K15n142188$ and $K_{-}=K15n156076$ are both given by gluing the exterior $E_{K3a1}$ of the trefoil knot $K3a1$ to a $1$-cusped hyperbolic manifold $M_{\pm}$. To finish the proof, it is enough to show that in both cases the manifold $M_{\pm}$ is not the exterior of a knot in $S^{3}$.
From Table 4 we see that $K_{\pm}$ are satellites of the trefoil $K3a1$ with pattern $\pm L6a1$. It follows that $M_{\pm}$ is given by $w_{\pm}$-surgery on one component [^3] Since $L6a1$ has an isometry that interchanges the components (and preserves the framing) it does not matter which component we fill. of $\pm L6a1$, where $w_{\pm}=12\pm 3$ is the writhe of the diagrams of $K_{\pm}$ shown in Figure 1. From that description, we can build the manifold $M_{\pm}$ in SnapPy, compute the short slopes (i.e. those of length at most $6$) of $M_{\pm}$, and verify that the corresponding fillings are not homology spheres. Then the $6$-theorem [5, 37] implies that $M_{\pm}$ is not the exterior of a knot in $S^{3}$.
For the census knots, all exceptional fillings are classified by Dunfield [22]. From his list, we can read off as above that no 0-surgery can be a connected sum. ∎
@@
$K14n22180$ (3)(-3)-fold twisted positive Whitehead double of $K3a1$
$K14n26039$ (3)(-3)-fold twisted negative Whitehead double of $K3a1$
@@
$K15n59184$ (2)(-2)-fold twisted negative Whitehead double of $K3a1$
$K15n115646$ (4)(-4)-fold twisted positive Whitehead double of $K3a1$
@@
$K15n142188$ (3)(-3)-fold twisted satellite of $K3a1$ with pattern $L6a1$, see Figure 1
$K15n156076$ (3)(-3)-fold twisted satellite of $K3a1$ with pattern L6a1-L6a1, see Figure 1
Figure 1. Diagrams of $K15n142188$ (left) and $K15n156076$ (right) showing their satellite structures on $K3a1$. Both diagrams were created with KLO [49].
@@ 4. Mirrors
  • The mirror $K_{m}$ of $K_{m}$ is isotopic to $K_{-m}$,
  • $K_{m}$ is not isotopic to its mirror $K_{m}$, but
@@
We will describe the infinite family via an RBG link construction as for example in [40, 46, 47, 39, 11]. For that we consider the $3$-component hyperbolic link $RBG$ with components $R$, $G$, and $B$ in Figure 2(i). By rotating that diagram around the $z$-axis, we see that there exists an involution $f\colon S^{3}\setminus RBG\rightarrow S^{3}\setminus RBG$, that maps $R$ to itself but interchanges $G$ and $B$.
@@
Now we see $RBG$ as a Kirby diagram by dotting $R$ and adding framings m1-m-1 and m1m-1 to $G$ and $B$, respectively. Note that $R$ is a meridian of both $G$ and $B$ (but not simultaneously). Thus we can cancel $R$ and $G$ and denote the image of $B$ under this cancellation by $K_{B}^{-m}$. By symmetry, we can also reverse the roles of $B$ and $G$ and get a knot $K_{G}^{m}$. By construction, these knots share the same trace. From the above symmetry of the RBG link, it follows that $K_{B}^{m}$ is isotopic to $K_{G}^{m}$ and we define this knot to be $K_{m}$. Thus we have constructed knots $K_{m}$ such that $K_{m}$ and $K_{-m}$ share the same trace. Figure 2(ii)-(iii) show explicit Kirby moves to create out of the RBG link the knots $K_{m}$.
@@
Next we show that $K_{m}=K_{-m}$ and that the $K_{m}$ are mutually distinct. Since the $K_{m}$ are obtained from $K_{0}$ by twisting three strands $m$ times, it follows from [34] that infinitely many of the $K_{m}$ are distinct. More concretely, we can see that they are all mutually distinct. For that, we consider the surgery description of the $K_{m}$ along the $2$-component link $L$ shown in Figure 2(iv). With SnapPy we check that $L$ is a hyperbolic link. By computing the volumes of the $K_{m}$ for small values of $m$ and using [42] we conclude that the volumes of the $K_{m}$ converge monotonically to the volume of $L$ and in particular all $K_{m}$ (for $m\in\mathbb{N}$) have different volumes.
@@
Moreover, we use SnapPy to compute the symmetry group of $L$, from which we observe that $L$ is isotopic to its mirror. This implies that $K_{-m}$ is isotopic to the mirror $K_{m}$ of $K_{m}$.
@@
It remains to demonstrate the each of the $K_{m}$ is not isotopic to its mirror. This can also be deduced from the symmetry group of $L$. Using SnapPy, we check that the complement of $L$ admits no non-trivial orientation-preserving isometry that extends to the link. Thus Thursthon’s hyperbolic Dehn filling theorem [53] implies that any sufficiently large Dehn filling of $L$ yields manifolds without any isometry, see [9] for details. The $K_{m}$ arise by Dehn filling one component of $L$ with slope $1/m$. Since the length of $1/m$ converges monotonically to $\infty$ it follows that infinitely many of the $K_{m}$ are asymmetric and thus not isotopic to their mirrors. More concretely, we can show that in fact for all $m\in\mathbb{N}$ the knots $K_{m}$ are not isotopic to their mirrors. For that we use the quantification of Thurston’s hyperbolic Dehn filling theorem [25] (see for example [9]) to compute an explicit bound $B$ such that $K_{m}$ is asymmetric for all $m\geq B$. For the finitely remaining $m$, we use SnapPy to create triangulations of the complements of $K_{m}$ and check that their symmetry groups do not contain any orientation-reversing diffeomorphisms. (But $K_{1}$ and $K_{2}$ turn out to be not asymmetric.) ∎
m1-m-1
m1-m-1
$K_{B}^{-m}=K_{-m}$
m1m-1
m1m-1
0
@@
Let $K$ be a prime knot with at most $15$ crossings or a hyperbolic knot with tetrahedral complexity at most $9$. Then $K$ and K-K have orientation-preserving diffeomorphic 0-surgeries if and only if $K$ is amphicheiral or if $K$ is isotopic to one of the following six chiral knots: $K13n469=v2272$, $K14n3411$, $K15n64176$, $t11462$, o9_22951o9\_22951, and o9_41058o9\_41058.
@@
If $K$ and $K^{\prime}$ have orientation-preserving diffeomorphic 0-surgeries then their signatures agree $\sigma(K)=\sigma(K^{\prime})$. For any knot $K$ we have $\sigma(K)=-\sigma(-K)$. Thus if $K$ and K-K have orientation-preserving diffeomorphic 0-surgeries, then the signature of $K$ has to vanish.